CARA Telecon Minutes Date: April 9, 2001 Time: 1:00 PM Attendance: BU: Cornell: CMU: Northwestern: G. Novak R.I.T: S.A.O.: T. Stark, A. Lane, W. Walsh South Pole: C. Martin, C. Kaminski Chicago: J. Carlstrom, S. Meyer, N. Odalen R. Pernic, R. Landsberg, J. Bausch, M. Thoma UCB: UCSB: J. Ruhl Approval of previous minutes: Motion made and seconded Agenda Items: News from the Pole: Moonrise was 2 days ago. There has been a medical problem at Pole this past week that has taken up most of the time at station. This problem has cleared, but may lead to an airdrop within 2 weeks. M. Newcomb is in need of parts for the Viper telescope. J. Carlstrom contacted D. Bresnanhan of NSF. CARA will be notified if an airdrop will take place. C. Martin reported that AST/RO has continued with observing the SMC in CO 2-1. This work is almost completed and then they will let the 1.3 mm receiver warm up and cool down Wanda to start 800 GHz observing. The first focus will be large scale mapping of the Galactic center region. Additional problems with Viper - 2 cables in the azimuth cable wrap were found to be cut when the wraps were inspected after the covers were taken off. The cut cables do not seem to effect anything on the telescope. M. Newcomb thinks that the cables maybe for CARONA or SPARO. M. Newcomb will fix the cable and he will take steps to make sure these cables are protected in the future. The cause was that the cable had too much slack; looped out of the wrap and got caught, maybe on the wheels? ***J. Ruhl to send G. Novak web address so that he can review the video that was taken by M. Newcomb of the cables. C. Kaminski commented that one cable was damaged, and the other cable looked OK. The cable was caught between the wheels and the cable wrap. They are unable to determine if the cable was being pulled. They are still pursuing with the problem solving. C. Kaminski has gotten one motor together. AST/RO - as reported by T. Stark. AST/RO is observing the SMC. The data looks good. AST/RO is observing the same area as the ISO satellite in CII. AST/RO is observing this area in CO 2-1. DASI - J. Carlstrom reported that the new pointing model is running and cured the tilt problem. Tests are being done on the effectiveness of the DASI ground shield for polarization observations - final observing strategy and 'relentless' observing should start very soon. The Chicago group of DASI has focused on getting last years data out. They are still looking at a completion time at the end of April'01. The power spectrum and cosmological parameter extraction is essentially done and the group is now writing. AASTO - C. Kaminski reported that the ice stack is about the same. Everything is being monitored and is working fine. C. Kaminski is physically going into the AAGO every 2 days to check on the TEG. Power dissipation is being monitored. Tipper on AST/RO is still working. Acbar - will be doing drift scans while the telescope is down. The drift scans (with the chopper running) will produce good CMB anisotropy data. The fields will also overlap with the DASI and Boomerang fields. This is likely to produce very good results and therefore there is less pressure to rush the debugging and fixing of the Viper azimuth cable wrap and drives. Kudos to C. Kaminski for helping M. Newcomb with the problems that have been creeping up on the telescope. Liquid Helium - C. Kaminski has sent J. Ruhl a spreadsheet on the status of the helium available at Pole. Westington #2 now has 700 liters available. The other Westington has 2500 liters and the 3000 gallon has had helium drawn but not too much. The boil off rates of the Westington is 20-25 liters. The thought was to transfer the 700 liters of helium out of the one Westington, but now this is almost a non-issue. They should transfer this helium to one of the 500 liter dewars. G. Novak feels comfortable with the information he reviewed on how the transfer was done at McMurdo to the 3000 gallon dewar. G. Novak believes that the 3000 gallon was filled at McMurdo. R. Pernic will be having a telecon with Raytheon on how the SIPs will be submitted this year. R. Pernic stated that conflicting information has been received from Raytheon and this telecon should clear up the procedures that will be followed by CARA. J. Carlstrom stated that he does not want SIPs submitted with separate SENs. All SIPs should be filed under SEN 370. *** a follow email explained the procedure which was deveoped in coordination with Ratheon*** *** J. Ruhl to coordinate on the helium request for the next season. ** a follow up email to Ruhl reviewed the plan to include both instrument usage and amount needed at station in CARA SIP ** J. Ruhl requested that each group report to both he and R. Pernic on usage for the upcoming season. NSF funds are in the 'process' of being electronically transmitted from NSF Grants and Agreements to the University of Chicago. N. Odalen will notify each group once the funds are received at Chicago. The UCSB & UCB financial people will be notified once these funds are at Chicago to allow spending by the PIs of these institutions. ***J. Carlstrom will call D. Peacock to see what is happening on the new proposals. R. Landsberg reported that 5 REUs have accepted. These individuals will need places to live during their term of employment. There are additional REU funds available. Each CARA group should review their needs and report to R. Landsberg. SIPs Equipment List - S. Meyer to give J. Carlstrom oscilloscope numbers (action item from last telecon). T. Stark requests that a RF meter and power head be included on this list. One was requested and received last year from Raytheon, but another is still needed.